40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 03:18:14 2011, kiwamu, Update, LSC, Y arm locking prep image.png
    Reply  Tue Sep 13 19:31:09 2011, kiwamu, Update, LSC, Y arm locking prep Yarmlocked.png
       Reply  Wed Sep 14 01:17:51 2011, kiwamu, Update, LSC, measurement of spot position on Y arm misposition.png
          Reply  Wed Sep 14 15:00:00 2011, kiwamu, Update, LSC, measurement of spot position on Y arm DSC_3408_small.jpg
    Reply  Wed Sep 14 21:18:45 2011, kiwamu, Update, LSC, Y arm locking prep 
Message ID: 5398     Entry time: Tue Sep 13 19:31:09 2011     In reply to: 5392     Reply to this: 5400
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Y arm locking prep 

The Y arm has been locked with AS55.

A next thing is to check the spot positions on the ETMY and ITMY mirrors so that we can evaluate the recent beam pointing.

 

- - - parameter settings - - -

C1:LSC-YARM_GAIN = -0.03

AS55 demod phase = 0.2

WF gains = 21 dB

C1:LSC-TRY_OUT = 0.57 (maximized by steering PZT2)

Quote from #5392

This was the reason why I had to do the f2a filter redesign. Tomorrow we can try locking the Y arm.

 

Attachment 1: Yarmlocked.png  27 kB  Uploaded Tue Sep 13 20:54:15 2011  | Hide | Hide all
Yarmlocked.png
ELOG V3.1.3-